# Fast LISA likelihood approximations by downsampling

### v1

@@ 4.4.2 Waveform model
The parameterised waveform model we use is the inspiral stage of the time-domain waveform of general relativistic binary black holes, as detailed in Ref. [25], with the following 8 free parameters: chirp mass ($\mathcal{M}_{\mathrm{c}}$), (inverted) mass ratio ($q$), effective spin ($\chi_{\mathrm{eff}}$), distance ($d$), inclination angle ($\theta_{\mathrm{JN}}$), polarisation angle ($\psi$), coalescence time ($t_{\mathrm{c}}$), and finally, the coalescence phase ($\phi_{\mathrm{c}}$), which, however, is marginalised numerically in the likelihood (see Appendix B), leaving 7 free parameters in the final posterior. To convert the waveform parameters into the signal property variables noted in the previous subsection, we can invert the well-known formulas for the inspiral part of BHB waveforms (see, e.g., Ref. [26])
@@
As mentioned in the previous subsection, changing the BHB parameters of course modify the posterior, but are less directly connected with the effects of downsampling. Hence the remaining signal model injection parameters for all fiducial systems will be made equal across the set of fiducial systems, and given in Table 4.4.2.
@@ 4.4.3 Detector model
In practice, the frequency of the measured waveform will be modulated by LISA’s motion around Sun. This effect makes sky localisation very precise and introduces sidebands into the waveform. In order to keep the frequency range small, in part to help to keep the MCS small and waveform evaluation costs down (we will be producing very many posteriors which can be expensive, so we take some measures to reduce the cost of this analysis), we omit the sky position waveform modification and fix the sky position. We also ignore the tumbling motion and frequency response of LISA, effectively modelling LISA as a LIGO-like detector in a plane with normal vector pointing directly at the source, with stationary noise.
@@ 4.4.4 PSD and MCS
In order to decrease the sampling rate, we can high-cut/low-pass the PSD at just above the Nyquist rate of the signal, since the frequency domain inner product [26] is unaffected if one takes the upper integration limit $\infty\to f_{\mathrm{max}}$, where
$f_{\mathrm{max}}=\max(\operatorname{supp}(\tilde{a}^{*}\tilde{b}))\,,$
and where $\operatorname{supp}(\cdot)$ stands for the support of the given function. However, what we are really interested in is minimising the MCS, i.e., the number of datapoints of the residual required to be computed, with minimal affect of the inner product. Simplifying the integral further, we can write
@@
where
$f_{\mathrm{min}}=\min(\operatorname{supp}(\tilde{a}^{*}\tilde{b}))\,,$
(note that we use a slightly larger $f_{\mathrm{max}}$ and slightly smaller $f_{\mathrm{min}}$ so as to account for the other template signals that will have a non-negligible probability in the posterior). Therefore, the behaviour of $S_{\mathrm{n}}(f)$ outside of the range $f_{\mathrm{min}}<f<f_{\mathrm{max}}$ is inconsequential to the inner product, and may be altered in any way necessary so as to minimise the MCS. We have not devised any such optimisation procedure, but have found that roughly flattening the PSD outside the signal frequency range, by simply using the new PSD given by
$S_{\mathrm{n}}^{\prime}(f)=\begin{cases}S_{\mathrm{n}}(f_{\mathrm{min}})&\text{for }f<f_{\mathrm{min}}\,,\\ S_{\mathrm{n}}(f)&\text{for }f_{\mathrm{min}}<f<f_{\mathrm{max}}\,,\\ S_{\mathrm{n}}(f_{\mathrm{max}})&\text{for }f>f_{\mathrm{max}}\,,\end{cases}$
can yield a low MCS; the MCS for most of our testbed systems (introduced in Section 4.6) with unaltered PSD is around 150, and drops to around 5 after this PSD modification; about 14 times fewer datapoints are required. Of all of our testbed systems after PSD modification, the highest MCS is 7, so we will override the MCS for the analyses of all other systems to be equal to 7, in order to be able to compare the results on posterior evaluation time like-for-like.
Finally, we note that we introduce another multiplicative factor to the PSD, $S_{\mathrm{n}}^{\prime}(f)$, such that the SNRs of all the injected waveform are equal to 8. The reason for this is that we are interested in accurately reproducing any given likelihood function; for higher SNRs, the likelihoods approach simple Gaussians (i.e., the lower the SNR, the more structure one can expect), and in GW data analysis there is a generally accepted SNR detection requirement of $\mathrm{SNR}\gtrsim 8$, i.e., the lower bound should produce the highest degree of posterior structure (least similar to a Gaussian), while still being detectable.
@@ 4.5 Posterior production
To acquire posterior estimates, we use the python Bayesian inference package bilby [27], to run the nested sampling sampler nessai [22]. The results are sets of points that are estimate draws from the posterior defined in terms of the likelihood function that we have introduced in this paper. The likelihood function that implements the downsampling procedures we have developed is provided by the python package dolfen, see Section 6 for further details.
@@ 4.6 The distances evaluated
4.6.1 Part 1: Best downsampling scheme and method
@@
To determine the best downsampling scheme and method, we use one combination of the signal injection parameters as described in Section 4.4.1, chosen to be the least amenable to downsampling. That is, that which is least well described on the whole as slowly evolving (the expected Fisher information changes most over the duration of the signal), and that which displays the highest degree of structure in the posterior. For our GW BHB model, this is where $N_{\mathrm{f}}=10^{6}$ and $\Delta f/f_{\mathrm{max}}=0.9$.
Figure 1: Convergence of the downsampling scheme (datapoint choice) and method (how to define the new inner product operator). This plot shows the mean of the posterior noise (the mean of the variation in the CMJS distances) induced by downsampling, for sets containing 21 posteriors defined using different random time series sample draws for various downsampling schemes and methods, for posterior sets defined by various $N_{\mathrm{s}}$. The best performing method and scheme under this metric is that which drops and stays below the (zeroed mean) physical noise (black dashed line) first (from the left). There is roughly equal performance between proposed schemes and methods with the exception of single factor cluster downsampling. The best performing scheme and method, though by a small amount, appears to be single factor hybrid downsampling.
The target distribution was defined with $\hat{N}_{\mathrm{s}}=32768$ and complied with the requirements set out in Section 4.3.4, where we chose $\hat{N}^{\prime}_{\mathrm{s}}=8192$. The downsampling scheme and method we used for the bootstrapping procedure was random sampling and minimisation of Jeffreys’ divergence of PCMs. We computed
  • $\mu[X_{\mathrm{smp}}]\approx 4.78\times 10^{-4}\,\mathrm{bits}$, and $\mu[X_{\mathrm{DS+smp}}]\approx 5.64\times 10^{-4}\,\mathrm{bits}$
  • $\sigma[X_{\mathrm{smp}}]\approx 1.33\times 10^{-4}\,\mathrm{bits}$, and $\sigma[X_{\mathrm{DS+smp}}]\approx 2.60\times 10^{-4}\,\mathrm{bits}$
  • $\mu[Y^{q}_{\mathrm{DS+smp}}(\hat{N}_{\mathrm{s}})]\approx 5.06\times 10^{-3}\,\mathrm{bits}$, and $\mu[Y^{q}_{\mathrm{DS+smp}}(\hat{N}^{\prime}_{\mathrm{s}})]\approx 6.67\times 10^{-3}\,\mathrm{bits}$
  • $\sigma[Y^{q}_{\mathrm{DS+smp}}(\hat{N}_{\mathrm{s}})]\approx 8.53\times 10^{-3}\,\mathrm{bits}$, and $\sigma[Y^{q}_{\mathrm{DS+smp}}(\hat{N}^{\prime}_{\mathrm{s}})]\approx 1.19\times 10^{-2}\,\mathrm{bits}$
@@
Thus we can see that the first set of conditions as summarised in Section 4.3.4 are satisfied. The best downsampling scheme and method is that which satisfies the second set of conditions given in Section 4.3.4. The results are displayed in Fig. 4.6.1, where we can see that the hybrid scheme using the method of minimisation of Jeffreys’ divergence of PCMs demonstrates the best performance.
4.6.2 Part 2: Optimum downsampling rate
@@
We shall now use the optimal downsampling method and scheme found in the previous subsection, and determine its downsampling posterior noise statistics for the set of fiducial systems given in Section 4.4.1. The results are shown in Figure 2.
@@
With a strict application of the convergence criteria summarised in Section 4.3.4, we find, from Figure 2, that the optimal number of datapoints depending on $\Delta f/f_{\mathrm{max}}$ appears to be as per Table 4.6.2.
$N_{\mathrm{s}}^{\mathrm{opt}}(\mathcal{H}_{q})$ $N_{\mathrm{f}}=10^{6}$ $N_{\mathrm{f}}=10^{7}$ $N_{\mathrm{f}}=10^{8}$
$\Delta f/f_{\mathrm{max}}=0.9\phantom{00}$ 362 (11.1) 362 (12.6) 362 (15.2)
$\Delta f/f_{\mathrm{max}}=0.09\phantom{0}$ 16 (12.7) 1448 (16.9) 181 (16.0)
$\Delta f/f_{\mathrm{max}}=0.009$ 1024 (10.0) 2048 (10.9) 16 (12.5)
Table 2: The optimal number of samples and the entropy of a representative target distribution, for all test systems. These values derive from Figure 2 and the convergence criteria given in Section 4.3. The green cells show expected behaviour and comply with convergence criteria in equation (4.3.3). The yellow cells required condition (4.3.3) for acceptance, and the red show unexpected behaviour. The cause of this behaviour appears to be from disproportionate information loss across the parameters, and the solution is to use FIM preservation as the downsampling method; this leads to vastly improved results (red dots/crosses in Figure 2), where $N_{\mathrm{s}}^{\mathrm{opt}}\leq 128$.
@@
Some surprising and interesting results are immediately apparent, highlighting important features of the analysis and limits of applicability of the downsampling method. These behaviours appear to result from a combination of the particulars of our signal model (how the data point values are derived from the parameters), and particulars of each posterior morphology (including the effects of the prior probability distributions on them).
Figure 2: The means and standard deviations of the sets of CMJS divergences, $\overline{D}_{\mathrm{JS}}$, as functions of $N_{\mathrm{s}}$, for the systems presented in Section 4.4.1. The top axis shows the fraction of the signal’s number of Nyquist samples that were required to be computed to evaluate the posteriors. The solid blue line (/blue crosses) shows the CMJS distance mean (/standard deviation) of sets of downsampled posteriors defined using single noise factor, hybrid downsampling, from the target distribution. The solid (/dashed) green lines are the means (/standard deviations) of the posterior noise of zeroed mean distributions defined with physical non-zero noise realisations. The red dots (/crosses) are the means (/standard deviations) of CMJS distances defined using the FIM preservation downsampling method. Convergence criteria is given in Section 4.3, and Table 4.6.2 gives $N_{\mathrm{s}}^{\mathrm{opt}}$ for each system.
@@
In Figure 2 and Table 4.6.2, we see that the behaviour of the downsampling for all $\Delta f/f_{\mathrm{max}}=0.9$ systems is as expected, and independent of $N_{\mathrm{f}}$. The other results are far more erratic with no discernible pattern, however, on inspection of the posteriors and consideration of the signal model and downsampling process, an explanation emerges and an important aspect of downsampling is highlighted.
On examining the FIMs of the posteriors that are very divergent from the target distributions, which cause the high degree of posterior noise, we find that they generally have in common that the information on the effective spin parameter far exceeds what should be expected, proportionately, with respect to the other parameters. Less generally, other effects are apparent on comparing the posteriors and the FIMs, such as relative information on certain pairs of parameters being disproportionately represented. Some of these subtle effects are compounded by the non-white noise covariance matrix and become very difficult to disentangle and investigate. However, since the FIM is the main moment of the distribution describing its morphology, ensuring the Fisher information is preserved after downsampling seems to be a possible approach to achieving a better downsampling rate for these extreme cases.
@@
To test this hypothesis, we computed specific sets of posteriors, 21 in each set, for those systems that converged much later than expected. We used the FIM preservation method rather than single factor noise reduction, since this method is of course intended to provide accurate covariances. The means and standard deviations of the particular sets we computed are shown as the red dots (means) and crosses (standard deviations) in Figure 2, and indeed very strongly support this hypothesis. Note that for $N_{\mathrm{f}}=10^{6},\Delta f/f_{\mathrm{max}}=0.009$ we used random rather than hybrid sampling, since dolfen had difficulty finding a solution for FIM preservation in the hybrid case.
@@
This appears to demonstrate that we initially chose a fiducial system with which to determine the optimal downsampling scheme and method in Part 1 that was sub-optimal, and which ultimately has led to the wrong choice of downsampling scheme with which to determine the optimal number of samples. Clearly using FIM preservation allows $N_{\mathrm{s}}^{\mathrm{opt}}\leq 128$ for those systems requiring $N_{\mathrm{s}}^{\mathrm{opt}}\geq 1024$ under single factor downsampling. Replacing the values in the red cells of Table 4.6.2 with the new values, $N_{\mathrm{s}}=128$, we can conclude that, using FIM preservation downsampling, the optimal downsampling number of samples for our range of tested systems is at most $N_{\mathrm{s}}^{\mathrm{opt}}=362$. In many cases it is much smaller than this, however, a method of computing $N_{\mathrm{s}}^{\mathrm{opt}}$ for a specific system (without producing and studying many posteriors as we have done) from, say, the Fisher matrix, presently does not appear to be possible.

### v2

@@ 4.4.2 Waveform model
The parameterised waveform model we use is the inspiral stage of the time-domain waveform of general relativistic binary black holes, as detailed in Ref. [IMRPhenomTP], with the following 8 free parameters: chirp mass ($\mathcal{M}_{\mathrm{c}}$), (inverted) mass ratio ($q$), effective spin ($\chi_{\mathrm{eff}}$), distance ($d$), inclination angle ($\theta_{\mathrm{JN}}$), polarisation angle ($\psi$), coalescence time ($t_{\mathrm{c}}$), and finally, the coalescence phase ($\phi_{\mathrm{c}}$), which, however, is marginalised numerically in the likelihood (see Appendix B), leaving 7 free parameters in the final posterior. To convert the waveform parameters into the signal property variables noted in the previous subsection, we can invert the well-known formulas for the inspiral part of BHB waveforms (see, e.g., Ref. [Maggiore])
@@
As mentioned in the previous subsection, changing the BHB parameters of course modifies the posterior, but is less directly connected to the effects of downsampling. Hence, for all fiducial systems we keep the remaining injection parameters fixed, as listed in Table 4.4.2.
@@ 4.4.3 Detector model
In practice, the frequency of the measured waveform will be modulated by LISA’s motion around the Sun. This effect enhances sky localisation and introduces sidebands into the waveform. To keep the frequency range small (and thereby reduce the MCS and waveform-evaluation cost for this large suite of posteriors), we omit the sky-position modulation and fix the sky location. We also ignore the tumbling motion and frequency response of LISA, effectively modelling LISA as a LIGO-like detector in a plane with normal vector pointing directly at the source, with stationary noise.
@@ 4.4.4 PSD and MCS
In order to decrease the sampling rate, we can high-cut/low-pass the PSD at just above the Nyquist rate of the signal, since the frequency-domain inner product [Maggiore] is unaffected if one takes the upper integration limit $\infty\to f_{\mathrm{max}}$, where f_max=max(supp(~a^*~b ))   , and where $\operatorname{supp}(\cdot)$ stands for the support of the given function. However, what we are really interested in is minimising the MCS, i.e., the number of data points of the residual required to be computed, with minimal affect of the inner product. Simplifying the integral further, we can write
@@
where f_min=min(supp(~a^*~b ))   , (note that we use a slightly larger $f_{\mathrm{max}}$ and slightly smaller $f_{\mathrm{min}}$ so as to account for the other template signals that will have a non-negligible probability in the posterior). Therefore, the behaviour of $S_{n}(f)$ outside the range $f_{\mathrm{min}}<f<f_{\mathrm{max}}$ is inconsequential to the inner product, and may be altered in any way convenient for minimising the MCS. We have not attempted a systematic optimisation, but we find that roughly flattening the PSD outside the signal frequency range, by using
[…]
already yields a very low MCS. The MCS for most of our testbed systems (introduced in Section 4.6) with unaltered PSD is around 77, and drops to around 5 after this PSD modification; i.e. the number of required samples is reduced by roughly a factor of 14. Of all our testbed systems after PSD modification, the highest MCS is 7, so for all subsequent systems we fix $\mathcal{M}=7$. This allows like-for-like comparison of posterior evaluation times across the parameter space.
Finally, we introduce an overall multiplicative factor in $S^{\prime}_{n}(f)$ such that the SNRs of all injected waveforms are equal to 8. We are interested in accurately reproducing generic likelihood functions; at higher SNRs the likelihoods become increasingly Gaussian (lower SNRs exhibit more structure), and in GW data analysis an SNR threshold of $\mathrm{SNR}\gtrsim 8$ is commonly used for detection. Choosing $\mathrm{SNR}=8$ therefore yields posteriors with non-trivial structure while remaining representative of detectable signals.
@@ 4.5 Posterior production
To obtain posterior estimates, we use the Python Bayesian inference package Bilby [Bilby] with the nested sampler Nessai [Nessai]. The results are sets of samples drawn from the posterior defined by the likelihood function introduced in this paper. The likelihood implementing the downsampling procedures is provided by the Python package Dolfen; see Sec. 6 for further details.
@@ 4.6 The distances evaluated
4.6.1 Optimal downsampling scheme and method
@@
To determine the best downsampling scheme and method, we use one combination of the signal injection parameters described in Sec. 4.4.1, chosen to be least amenable to downsampling: namely, the system that is least well described as slowly evolving (its Fisher information varies most over the signal duration) and whose posterior is most highly structured. For our GW BHB model this corresponds to $N_{\mathrm{f}}=10^{6}$ and $\Delta f/f_{\mathrm{max}}=0.9$.
The target distribution was defined with $\hat{N}_{\mathrm{s}}=32768$ and satisfies the requirements in Sec. 4.3.4, with $\hat{N}^{\prime}_{\mathrm{s}}=8192$ used as the comparison point. For the bootstrapping procedure we use random sampling and the single–factor method based on minimising the Jeffreys’ divergence between parameter covariance matrices. Quantitatively, the CMJS divergences arising from sampler noise alone are of order $\sim 5\times 10^{-4}\,\mathrm{bits}$, and including downsampling increases this only slightly (to $\sim 6\times 10^{-4}\,\mathrm{bits}$). The CMKL divergences to the target at $\hat{N}_{\mathrm{s}}$ and $\hat{N}^{\prime}_{\mathrm{s}}$ are $\sim 5$$7\times 10^{-3}\,\mathrm{bits}$ with standard deviations of order $10^{-2}\,\mathrm{bits}$, so the additional variability from downsampling at these rates is small compared to the intrinsic scatter.
@@
Thus we can see that the first set of conditions as summarised in Sec. 4.3.4 are satisfied. The optimal downsampling scheme and method are those that satisfy the second set of conditions in Sec. 4.3.4. The corresponding results are shown in Fig. 4.6.1; the hybrid scheme combined with the single–factor method (minimising the Jeffreys’ divergence of PCMs) appears to yield the best performance.
Figure 1: Convergence of the downsampling scheme (data point choice) and method (how to define the new inner product operator). This plot shows the mean of the posterior noise (the mean of the variation in the CMJS distances) induced by downsampling, for sets containing 21 posteriors defined using different random time series sample draws for various downsampling schemes and methods, for posterior sets defined by various $N_{\mathrm{s}}$. The best performing method and scheme under this metric is that which drops and stays below the (zeroed mean) physical noise (black dashed line) first (from the left). There is roughly equal performance between proposed schemes and methods with the exception of single factor cluster downsampling. The best performing scheme and method, though by a small amount, appears to be hybrid downsampling combined with the single–factor Jeffreys–divergence method.
4.6.2 Optimal downsampling rate
@@
We proceed with the optimal downsampling method and scheme found in the previous subsection, and determine its downsampling posterior noise statistics for the set of fiducial systems given in Section 4.4.1. The results are shown in Figure 2.
@@
With a strict application of the convergence criteria summarised in Section 4.3.4, we obtain, from Figure 2, the optimal number of data points as a function of $\Delta f/f_{\mathrm{max}}$; these values are summarised in Table 4.6.2.
@@
Several interesting features are immediately apparent, highlighting both the strengths and the limits of the downsampling method. The observed behaviour seems to arise from a combination of (i) details of our signal model (how data-point values are determined by the parameters) and (ii) the geometry of each posterior (including the influence of the prior distributions).
@@
From Figure 2 and Table 4.6.2 we see that the behaviour of the downsampling for the $\Delta f/f_{\mathrm{max}}=0.09$, $N_{\mathrm{f}}=10^{8}$ system, and for all $\Delta f/f_{\mathrm{max}}=0.9$ systems, is as expected. The remaining cases show no obvious pattern. However, on closer inspection of the posteriors and their FIMs, and in light of the signal model and the downsampling procedure, an explanation emerges and an important aspect of downsampling becomes clear.
$\Delta f/f_{\mathrm{max}}$ $N_{\mathrm{f}}=10^{6}$ $N_{\mathrm{f}}=10^{7}$ $N_{\mathrm{f}}=10^{8}$
$0.9$ 362 (11.1) [P] 362 (12.6) [P] 362 (15.2) [P]
$0.09$ 16 (12.7) [C2] 1448 (16.9) [F] 181 (16.0) [P]
$0.009$ 1024 (10.0) [F] 2048 (10.9) [F] 16 (12.5) [C2]
Table 2: The optimal number of samples and the entropy (round brackets) of a representative target distribution, for all test systems. These values derive from Figure 2 and the convergence criteria given in Section 4.3. The cells marked [P] show expected behaviour and comply with convergence criteria in equation (29). The cells marked [C2] required condition (30) for acceptance, and those marked [F] show unexpected behaviour, the cause of which appears to be from disproportionate information loss across the parameters. The solution is to use FIM preservation as the downsampling method; this leads to vastly improved results (red dots/crosses in Figure 2), where $N_{\mathrm{s}}^{\mathrm{opt}}\leq 128$.
Examining the FIMs for posteriors that deviate strongly from the target distributions (and hence exhibit a high level of posterior noise), we find that they typically share the feature that the information on the effective spin parameter is disproportionately large compared with that on the other parameters. In other cases, the relative information content of specific parameter pairs is also anomalously large. Some of these subtler effects are further compounded by the non-white noise covariance matrix, making them difficult to disentangle. Since the FIM is the leading descriptor of the local geometry of the posterior, preserving the Fisher information under downsampling appears to be a natural and promising route to achieving better downsampling rates in these extreme cases.
@@
To test this hypothesis, we computed specific sets of posteriors, 21 in each set, for those systems that converged much later than the ‘well-behaved’ systems, using the FIM preservation method rather than single factor noise reduction, since this method is of course intended to provide accurate covariances. The means and standard deviations of the particular sets we computed are shown as the red dots (means) and crosses (standard deviations) in Figure 2, and indeed very strongly support this hypothesis. Note that for $N_{\mathrm{f}}=10^{6},\Delta f/f_{\mathrm{max}}=0.009$ we used random rather than hybrid sampling, since Dolfen had difficulty finding a solution for FIM preservation in the hybrid case.
@@
This demonstrates that we chose a fiducial system with which to determine the optimal downsampling scheme and method in Part 1 that was sub-optimal, and which ultimately has led to the sub-optimal choice of downsampling scheme, with which the bulk analysis of determining the optimal number of samples was performed. Clearly using FIM preservation allows $N_{\mathrm{s}}^{\mathrm{opt}}\leq 128$ for those systems requiring $N_{\mathrm{s}}^{\mathrm{opt}}\geq 1024$ under single factor downsampling. By replacing the values in the cells of Table 4.6.2 marked [F] with the new values, $N_{\mathrm{s}}=128$, we can still confidently conclude that using FIM preservation downsampling, the optimal downsampling number of samples for our range of tested systems is at most $N_{\mathrm{s}}^{\mathrm{opt}}=362$. Most often, one requires far fewer samples than this, however, a general method of computing $N_{\mathrm{s}}^{\mathrm{opt}}$ for any given system (without producing and studying many posteriors as we have done) from, say, the Fisher matrix, presently does not appear to be possible.
Figure 2: The means and standard deviations of the sets of combined marginal Jenson-Shannon (CMJS) divergences, $\overline{D}_{\mathrm{JS}}$, as functions of $N_{\mathrm{s}}$, for the systems presented in Section 4.4.1. The top axis shows the fraction of the signal’s number of Nyquist samples that were required to be computed to evaluate the posteriors. The solid blue line (/blue crosses) shows the CMJS distance mean (/standard deviation) of sets of downsampled posteriors defined using single noise factor, hybrid downsampling, from the target distribution. The solid (/dashed) green lines are the means (/standard deviations) of the posterior noise of zeroed mean distributions defined with physical non-zero noise realisations. The red dots (/crosses) are the means (/standard deviations) of CMJS distances defined using the FIM preservation downsampling method. Convergence criteria is given in Section 4.3, and Table 4.6.2 gives $N_{\mathrm{s}}^{\mathrm{opt}}$ for each system.